The formula below gives the cost (in £) for the car tom is hiring for his wedding.
Cost = 150 + 45 (h - 1)
h is the number of hours for which the car is hired.
How much will it cost Tom to hire the car for 8 hours.
I just did 195x7 =1365 which is wrong please explain how to work out
Thanks

Found 2 solutions by jim_thompson5910, ewatrrr:
Answer by jim_thompson5910(35256)

Cost = 150 + 45 (h - 1)

Cost = 150 + 45 (8 - 1) ... Plug in h = 8

Cost = 150 + 45 (7)

Cost = 150 + 315

Cost = 465

So it costs 465 pounds

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------

Answer by ewatrrr(24785)

 
Hi,
Cost = 150 + 45(h - 1)
h is the number of hours for which the car is hired.
How much will it cost Tom to hire the car for 8 hours.
Cost = 150 + 45·(8-1)
Cost = 150 + 45·7 || Need to first determine 45(8-1) and then add to 150